OpenStudy (toxicsugar22):

The mass ratio of sodium to fluorine in sodium fluoride is 1.21:1. A sample of sodium fluoride produced 32.5 \({\rm g}\) of sodium upon decomposition. How much fluorine (in grams) was formed?

OpenStudy (aaronq):

\(\dfrac{Na}{F}=\dfrac{1.21}{1}\rightarrow\dfrac{1.21}{1}=\dfrac{32.5 g}{F}\)

OpenStudy (aaronq):

just gotta write what they told you in terms of a ratio

OpenStudy (toxicsugar22):

ok i got it itis 26.9 the answer right

OpenStudy (aaronq):

yes

OpenStudy (toxicsugar22):

can you help me with a couple more

OpenStudy (aaronq):

i have to go, sorry.